OpenStudy (needhelp):

for the log quetsion. ln 1= 0 <-in this case, do i need to assume e is there? i can't remember the formulas, is there any way i can remember it?

OpenStudy (anonymous):

ln 1 means to what power must you raise e to get 1.

OpenStudy (anonymous):

does that make sense?

OpenStudy (needhelp):

yea, I guess i am misunderstanding form of questions. can you help me out with more please?

OpenStudy (anonymous):

what are you having difficulty with?

OpenStudy (needhelp):

ln e = 1. i am confuse with the formulas

OpenStudy (needhelp):

what is the 1 stands for? is power?

OpenStudy (anonymous):

yes, in that case it is the power of e.

OpenStudy (anonymous):

i.e. ln e ^10 = 10.

OpenStudy (anonymous):

ln and e are inverses of each other.

OpenStudy (anonymous):

does that make sense?

OpenStudy (needhelp):

yea, so that means above that example i show, ln e^1, is this right?

OpenStudy (anonymous):

for ln e = 1 is just saying that the log of the base itself, which is e, is 1.

OpenStudy (anonymous):

Exponential to Logarithmic: \[b^{x} = a\] I use b because b is the Base. B is for Base. \[\log_{b} (a) = x\] Fill in the blanks to change it to logarithmic using the numbers from the exponential form. Always remember, A Log Is an Exponent. A log with Base "e" is a special log which isnt written as \[\log_{e} a\] , but rather it is written as \[\ln a\]

OpenStudy (needhelp):

so in this case, In a, means log a without the base?

OpenStudy (anonymous):

\[\ln a \] means \[\log_{e} a\] , but it just is written as \[\ln a \]

OpenStudy (needhelp):

so for example, if ln 9, what is the answer going to be?

OpenStudy (anonymous):

Therefore when you have ln e = 1 you actually have: \[\log_{e} e = 1\] Which, if you convert it back to exponential form, you get: \[e ^{1} = e\]

OpenStudy (anonymous):

ln 9 = \[\log_{e} 9\] which can convert to \[e ^{x} = 9\]

OpenStudy (needhelp):

for the exponent, is it possible to become a # instead of x ?

OpenStudy (anonymous):

ln 1=0 1 = e^0